C=100+0.50Y, Ip=100-20r, Mt=0.10Y, Ms=100-10r, M=80.

a. The IS–LM functions show the relationship between interest rates and real output, in the goods and services market and the money market.

The IS curve is defined by the equation Y = C + I + G + NX, so Y = 100 + 0.5Y + 100 - 20r - 80, Y = 240 - 40r. The LM curve is defined by the equation Mt = Ms, so 100 - 10r = 0.1Y, Y = 1000 - 100r.

b. If we assume an increase in Investments by 100 units, then the IS function will shift to the right.

c. The intersection of IS-LM functions defines four areas, and the behavior of the markets for goods and money for each area will change.


1
Expert's answer
2021-03-22T08:06:25-0400

a. The IS–LM functions show the relationship between interest rates and real output in the goods and services market and the money market.

The IS curve is defined by the equation Y = C + I + G + NX, as we have no government purchases G and exports X, then:

Y = 100 + 0.5Y + 100 - 20r - 80,

Y = 240 - 40r.

The LM curve is defined by the equation Mt = Ms, so:

100 - 10r = 0.1Y,

Y = 1000 - 100r.

b. If we assume an increase in Investments by 100 units, then the IS curve will shift to the right, both equilibrium interest rate and output will increase.

c. The intersection of IS-LM functions defines four areas, and the behavior of the markets for goods and money for each area will change in every area.